Radioactive decay of radioactive materials can be modeled by the equation


A=Aoe^t(in 2/h) , where A is the amount of the material at time t, A0 is the amount at time 0, and h is the half-life. Technetium-99 is a radioisotope that is used in imaging of the brain. It has a half-life of 6 hours. Your program should display the relative amount A/A0 in a patient body every hour for 24 hours after receiving a dose. Note: With the aid of formatting techniques documented in your note, ensure your program output matches that given below. Program Output


1: 0.890899


2: 0.793701


3: 0.707107


4: 0.629961


5: 0.561231


6: 0.500000


7: 0.445449


8: 0.396850


9: 0.353553


10: 0.314980


11: 0.280616


12: 0.250000


13: 0.222725


14: 0.198425


15: 0.176777


16: 0.157490


17: 0.140308


18: 0.125000


19: 0.111362


20: 0.099213


21: 0.088388


22: 0.078745


23: 0.070154


24: 0.062500

1
Expert's answer
2021-11-19T17:07:53-0500
from math import e
import numpy

A0 = int(input("A0:" ))
h = 6
t = 0
i = 0
print("Hours", " Millicuries")
while t < 25:
  final = (A0 * e **(-t*(numpy.log(2)/h)))/A0
  t = t + 1
  print(i, "    ", final)
  i = i + 1
#By Jasur007